OpenStudy (anonymous):

A ball has a diameter of 9 in. It consists of 2 parts. The inside is a spherical core with a diameter of 6 in. Surrounding the core is a layer of polyurethane. What is the volume of the polyurethane? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (freckles):

you will have to find the difference of the volume of the whole ball minus the volume of just the core to find the outer layer volume

OpenStudy (freckles):

use the sphere volume twice

OpenStudy (anonymous):

oh ok

OpenStudy (freckles):

try this one and put your answer here if it isn't what I have we will go over this slowly like the last

OpenStudy (anonymous):

radius 6 i got 113.04

OpenStudy (anonymous):

i mean diameter 6

OpenStudy (freckles):

wait

OpenStudy (freckles):

radius is not 6

OpenStudy (freckles):

diameter is 9 radius is 9/2

OpenStudy (anonymous):

i know it is 3

OpenStudy (freckles):

well for the core 3 is right for the radius

OpenStudy (anonymous):

i got for diameter 9 the radius 4.5 and got the volume of 381.51cm

OpenStudy (anonymous):

@freckles are you there

OpenStudy (anonymous):

so what do i do now

OpenStudy (freckles):

omg i keep writting 2 instead of 3

OpenStudy (anonymous):

its alright

OpenStudy (anonymous):

i dont get the last part in the bottom

OpenStudy (freckles):

you still need to find the volume of the smaller ball inside the big ball

OpenStudy (anonymous):

do i subtract the bigger number from the smaller number

OpenStudy (freckles):

the smaller from the bigger

OpenStudy (freckles):

doing the opposite will give you a negative number

OpenStudy (anonymous):

oh ok so 113.04 - 381.51

OpenStudy (freckles):

no take the smaller from the bigger lol \[\frac{4}{3} \cdot 3.14 \cdot (\frac{9}{2})^3-\frac{4}{3} \cdot 3.14 \cdot (\frac{6}{2})^3\\ 381.51-\frac{4}{3} \cdot 3.14 \cdot (\frac{6}{2})^3 \\ 381.51 -113.04\]

OpenStudy (anonymous):

sorry i meant the other way around

OpenStudy (anonymous):

268.47cm is what i got

OpenStudy (freckles):

great job except the units part cm^3 :)

OpenStudy (anonymous):

ok one more please

OpenStudy (freckles):

one more will have to wait after i run a quick errand

OpenStudy (freckles):

it will be kinda quick

OpenStudy (anonymous):

ok then thank you for your help anyway

OpenStudy (freckles):

i'm back

OpenStudy (anonymous):

ok do i start a new one or right here

OpenStudy (freckles):

I don't care one last question right?

OpenStudy (freckles):

if it is one question I will work on it with you

OpenStudy (anonymous):

ya last question

OpenStudy (anonymous):

A can of spaghetti has a diameter of 12 cm and a height of 16 cm. The can contains spherical meatballs, each with a radius of 3 cm, that take up 25% of the volume. How many meatballs are in the can? Use 3.14 to approximate pi and express your answer as a whole number.

OpenStudy (freckles):

now i'm hungry

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

lol

OpenStudy (freckles):

the spaghetti made me hungry :p

OpenStudy (freckles):

anyways

OpenStudy (freckles):

ok let's go ahead and find the volume of the can

OpenStudy (anonymous):

its a sphere volume right

OpenStudy (freckles):

well a can is a cyclinder

OpenStudy (anonymous):

oh ya my bad

OpenStudy (freckles):

we will use sphere formula for the meatballs though

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 1809.56

OpenStudy (freckles):

one sec checking

OpenStudy (freckles):

what did you do to get that (it is close to what I have but not exact)

OpenStudy (anonymous):

ya what did you get

OpenStudy (freckles):

1808.64

OpenStudy (freckles):

see kinda close

OpenStudy (anonymous):

i did it on google

OpenStudy (anonymous):

i just put volume of a cylinder and type in the radius and height

OpenStudy (freckles):

oh i see what you did

OpenStudy (freckles):

you put pi instead of 3.14

OpenStudy (freckles):

and it wants us to use 3.14 instead of pi but in reality your answer is more closer than mine

OpenStudy (freckles):

but we should do what the instructions want

OpenStudy (anonymous):

ok so your right. right

OpenStudy (anonymous):

1808.64

OpenStudy (freckles):

yah

OpenStudy (anonymous):

ok so what now

OpenStudy (freckles):

now try to find the volume of a single meatball :)

OpenStudy (anonymous):

what formula do i use

OpenStudy (freckles):

they describe meatballs as if they were spheres so you used the formula for volume of a sphere

OpenStudy (anonymous):

so it is 113.1

OpenStudy (freckles):

113.04 is what I have

OpenStudy (anonymous):

what so whos right

OpenStudy (freckles):

so pretend there are n meatballs so the volume that the meatballs make up is n*113.04 and we are given that n*113.04 is .25 of the volume we found from the cylinder that is: \[n \cdot 113.05=.25 \cdot 1808.64\] now we solve for n

OpenStudy (freckles):

\[n \cdot 113.04=.25 \cdot 1808.64\]*

OpenStudy (freckles):

well I did 4/3*3.14*3^3

OpenStudy (freckles):

what did you do?

OpenStudy (anonymous):

same thing. on the step you told me to do i dont get it

OpenStudy (freckles):

the equation involving solve for n?

OpenStudy (freckles):

we are given that a number of meatballs is 25% of the can's volume

OpenStudy (freckles):

I just translated that into a mathematical equation

OpenStudy (freckles):

i called the number n n*volume of meatball=.25*volume of can

OpenStudy (freckles):

divide both sides by volume of meatball to solve for n

OpenStudy (anonymous):

so whats the equation again

OpenStudy (freckles):

\[n=\frac{.25 \cdot 1808.64}{113.04}\]

OpenStudy (anonymous):

i got 4

OpenStudy (freckles):

sounds great

OpenStudy (freckles):

that sounds like a horrible spaghetti an because I prefer a lot of meatballs

OpenStudy (anonymous):

Thank you for all your help.

OpenStudy (freckles):

can*

OpenStudy (anonymous):

the answer is 4 right

OpenStudy (freckles):

yes that is what we came up with there are 4 meatballs

OpenStudy (anonymous):

hello